# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v2

@@ 5.1 A solution that naturally prevents overconfidence
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure 3). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure 1 (leftmost). Indeed, $F(N,\hat{p}(y|x))$ is smaller for easier examples (Figure 15, right), effectively regularizing their contribution. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section 4.4.
@@
We note in practice, the introduction of $F$ can lead to some samples in a batch contributing minimally to the current gradient step, thereby reducing the effective batch size. To maintain a stable effective batch size, we introduce a threshold $\epsilon$, and if for a given example, $(x,y)$, $F(N,\hat{p}(y|x))<\epsilon$, we replace the example with a new one.
Figure 3: The DCO objective limits overconfidence by attenuating gradients for examples on which the model is highly confident. We plot the confidence regularization factor $F(N,\hat{p}(y|x))$ in Equation 4 which attenuates CE loss gradients to obtain DCO loss gradients. $N=1$ corresponds to CE loss, where $F=1$ regardless of $\hat{p}(y|x)$. For any $N>1$, the attenuation is stronger ($F$ is smaller) for examples with higher model confidence $\hat{p}(y|x)$, and this attentuation effect increases with $N$. $F$ drops to near zero as soon as confidence $\hat{p}(y|x)$ increases to $1/N$. Thus as soon as an example’s confidence rises to $O(1/N)$ over training time, its gradient under the DCO loss cannot make the confidence rise any further. Thus DCO naturally limits overconfidence.
@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
We next test whether DCO can rescue test-time scaling by preventing model overconfidence. We first perform experiments on the MATH dataset in this section, and then in the next section we perform experiments on the LeanDojo automated theorem proving benchmark (Yang et al., 2023).
We fine-tune the LLama-3-8B-base model for 4 epochs on the MATH training set using DCO for $N=256$ and confirm that the model is far less confident on its greedy completion than when trained with CE loss after multiple training epochs (compare Figure 1 right and Figure 1 leftmost). Moreover, training with DCO at larger $N$ yields lower model greedy confidences at the end of training (Figure 1 rightmost), consistent with Lemma 4.2.
We next assess pass@N test performance as a function of $N$ for a variety of models trained by minimizing $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for different values of $N^{\prime}$ (Figure 4). For any given $N$ in pass@N, there is an optimal $N^{\prime}$ for the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ that maximizes pass@N test coverage, yielding a Pareto optimal performance frontier (black curve) that is achieved when $N^{\prime}$ is close to $N$. In particular the model trained with CE loss (equivalent to pass@1 maximization at training) performs poorly relative to the Pareto frontier at large N when the pass@N strategy is used at test time (red curve below black at large $N$). Conversely, models trained with DCO at large $N^{\prime}$ perform poorly relative to the Pareto frontier when the pass@N strategy is used at test time with small $N$ (green curves below black at small $N$).
Together these results indicate that the alignment of the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ with the test time strategy pass@N, with $N^{\prime}$ close to $N$, is crucial for obtaining Pareto optimal performance. Moreover, these results once again confirm the tradeoff between exploration and exploitation, with good performance using pass@N test-time strategies requiring high (low) exploration with low (high) confidence at large (small) $N$. Indeed, this tradeoff prevents achieving Pareto optimality using pass@N at test-time for all $N$ via training with DCO at any single value of $N^{\prime}$.
Figure 4: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$. (See Figure 8 for an example where the model trained with $N^{\prime}=256$ is highlighted.) Please refer to Figures 10 and 12 for corresponding results with the Llama-3-70B-base model and on the AIME24 dataset, respectively.
@@ 5.3 Improved theorem proving via ensembled tree search through a modified step-wise DCO
To further test our method in realistic settings with a verifier, we conduct experiments in theorem proving using an interactive proof assistant on the LeanDojo benchmark (Yang et al., 2023) extracted from the math library of LEAN4 (mathlib Community, 2020). In this task, at each step $i$ of the proof, the model is prompted with the current proof state $x[i]$, and it outputs a proof tactic $y[i]$ with a trainable model probability $\hat{p}(y[i]|x[i])$. The proof assistant takes the sampled tactic $y[i]$, verifies whether it is a valid tactic, and if so, returns the next proof state $x[i+1]$ after the tactic $y[i]$ is applied. The model then samples the next tactic $y[i+1]$. At test time this interactive process between the model and the assistant continues until either: (1) it terminates in an invalid tactic; (2) hits a maximal allowed search depth set by computational constraints; or (3) terminates in a successful proof of the initial proof goal as verified by the proof assistant.
@@
In our experiments, we use LEAN4 (Moura & Ullrich, 2021) as the formal proof language. We start from the model Qwen2.5-Math-1.5B (Yang et al., 2024), and fine-tune it on the LeanDojo benchmark (Yang et al., 2023). In contrast to solving math problems above where the model first autoregressively generates an entire answer $y$ according to $\hat{p}(y|x)$ and then the entire $y$ is verified for correctness, in theorem proving we must obtain correct tactics $y[i]$, as verified by the proof assistant, at every proof step $i$. A straightforward application of DCO fine-tuning in this setting then involves replacing the model confidence $\hat{p}(y|x)$ in  Equation 3 on a single math answer $y$, with the model confidence on an entire successful, complete $k$ step proof $\hat{p}(y[0],...,y[k-1]|x[0])=\prod_{i=0}^{k-1}\hat{p}(y[i]|x[i])$. Because of the chain rule, the DCO gradient on a single proof has the same form as in Equation 4 but with $F(N,\hat{p}(y|x))$ replaced with $F(N,\prod_{i=0}^{k-1}\hat{p}\left(y[i]|x[i]\right))$.
We naively applied this DCO fine-tuning method with $N=4$k and we evaluated the resulting model, as well as a baseline model trained with CE loss, on MiniF2F. As a test strategy, we used pass@4k. The baseline model fine-tuned with CE loss achieves a proof success rate of 37.4%, while model fine-tuned with DCO at $N=4$k achieves a 38.7% success rate. Thus, a naive application of DCO to theorem proving by matching the parameter $N$ in DCO to the parameter $N$ in pass@N achieves only a modest improvement.
@@
However, in theorem proving, since every single intermediate proof step tactic $y[i]$ must be valid, it is natural to consider a step-wise generalization of DCO. For example, at each step $i$, the model chooses a tactic $y[i]$ with confidence $\hat{p}(y[i]|x[i])$. Our proposed step-wise DCO optimizes this single-step confidence according to Equation 4 as before, except now the step-wise confidence regularizer is given by $F(N_{\text{eff}},\hat{p}(y[i]|x[i]))$. Here one can think of $N_{\text{eff}}$ as a step-wise DCO hyperparameter that controls the exploration width at every proof step. In essence, the confidence regularizer prevents the confidence of any chosen tactic from becoming much higher than $1/N_{\text{eff}}$. Since the sum of the confidences over all tactics at each step must be $1$, this means that at test time, model exploration at each step corresponds to searching on a search tree in which each proof state $x[i]$ allows the exploration of approximately only $N_{\text{eff}}$ tactics. Thus using $N_{\text{eff}}$ in step-wise DCO during fine-tuning selects the approximate branching factor $N_{\text{eff}}$ of the model’s proof search tree at test time (see  Figure 5).
Figure 5: Inverse confidence $N_{\text{eff}}$ at training time controls search-tree exploration shapes at test-time. We show, qualitatively, two proof search trees (solid lines) obtained when training at small $N_{\mathrm{eff}}$ (left) and large $N_{\mathrm{eff}}$ (right). Both trees start from the same theorem statement (top), circles denote proof states, directed edges denote tactics, solid edges are assigned non-negligible model confidence and are explored, while dashed edges are not. Small $N_{\mathrm{eff}}$ leads to a small branching factor (left), which can fail if the correct proof (red path) is not in the narrow search tree. Larger $N_{\mathrm{eff}}$ expands the search tree to include the correct proof (right). However, if $N_{\mathrm{eff}}$ is too large ($N_{\mathrm{eff}}^{k}\gg N$ for a theorem with $k$ proof steps), the sampling probability of the correct proof will decrease as $N_{\mathrm{eff}}^{-k}$ and the pass@N coverage $\mathcal{C}^{N}\approx N/N_{\mathrm{eff}}^{k}$ decreases as $N_{\mathrm{eff}}$ increases. In essence the correct proof is too hard to find in a search tree that is too wide. Overall, we see that inverse confidence $N_{\mathrm{eff}}$ chosen at training time constitutes a simple but powerful knob to control a tradeoff between exploitation and exploration on proof search trees of varying shapes at test time.
@@
In particular, a valid (partial) proof of length $k$ will have probability of order $N_{\text{eff}}^{-k}$. Thus small $N_{\text{eff}}$ limits the exploration at test time to a tree with small branching factor, but allows longer proofs with larger numbers of steps $k$ to have higher sampling probability. This limited width search strategy should work well at test time using pass@N as long as two conditions hold: (1) a successful proof of length $k$ is present in the search tree of branching factor $N_{\text{eff}}$; and (2) the $N$ in pass@N at test time is large enough that this proof of probability $O(N_{\text{eff}}^{-k})$ is selected with high probability in $N$ passes (which starts to occur as soon as $N\gtrsim N_{\text{eff}}^{k}$).
Conversely, larger $N_{\text{eff}}$ allows for more exploration at test time using a wider search tree of larger branching factor, but it makes finding longer proofs with large $k$ harder, since the approximate success condition for pass@N of $N>N_{\text{eff}}^{k}$ is harder to satisfy at fixed $N$ and large $N_{\text{eff}}$ and $k$. However, this wide exploration strategy can work well for short proofs with small $k$ if a successful short proof does not lie in the limited width search tree obtained at small $N_{\text{eff}}$, but does lie in a wider search tree at larger $N_{\text{eff}}$ (Figure 5, right).
We note that the mean and median lengths of the proofs in the training set is 4.2 and 2 respectively. Thus we explore fine-tuning on LeanDojo with the step-wise DCO algorithm DCOstep with $N_{\text{eff}}$ ranging from $1$ (corresponding to standard CE-loss training) to $N_{\text{eff}}=32$. We evaluated the pass@N test performance for $N=4$k of each of the resulting models on Mathlib and MiniF2F test sets, finding improved performance at larger $N_{\text{eff}}$ compared to the baseline $N_{\text{eff}}=1$ CE loss training (Table 2). Interestingly, we find that the optimal $N_{\text{eff}}$ increases with more passes at test time (Table 6), similar to the Pareto frontier in Figure 4.
@@
Since different choices of $N_{\text{eff}}$ correspond to very different search strategies at test time, specializing in finding short proofs on wide search trees for large $N_{\text{eff}}$ and long proofs on narrow search trees for small $N_{\text{eff}}$, we hypothesized that ensembling these methods could significantly boost performance. We found this was indeed the case (ensemble entry in Table 2). Of course the Ensemble strategy samples $5\times 4$k times, so a proper baseline is CE loss with the same test-time compute of pass@N with $N=20$k. We obtained the proof success rate for this baseline to be 57.0% on Mathlib and 39.5% on MiniF2F. The ensemble strategy outperforms this stringent baseline with significant excesses of $5.2\%$ on Mathlib and $4.1\%$ on MiniF2F.
@@
Overall, these results indicate that varying $N_{\text{eff}}$ in step-wise DCO at training time allows a diversity of tree search strategies trading depth and breadth that can be effectively exploited by simply scaling test time compute via pass@N.
Table 2: Proof success rate using pass@4k at test time on Mathlib and MiniF2F and training with DCOstep for different $N_{\text{eff}}$.
Mathlib MiniF2F
$N_{\mathrm{eff}}=1$ (CE loss) 55.6% 37.4%
$N_{\mathrm{eff}}=4$ 56.4% 39.0%
$N_{\mathrm{eff}}=8$ 56.1% 39.5%
$N_{\mathrm{eff}}=16$ 56.5% 37.0%
$N_{\mathrm{eff}}=32$ 55.8% 37.4%
Ensemble of above $N_{\mathrm{eff}}$ 62.2% 43.6%
$N_{\mathrm{eff}}=1$ (5x test compute) 57.0% 39.5%
@@
5.4 Approximate DCO also improves MATH answering with chain-of-thought reasoning

### v3

@@ 5.1 A solution that naturally prevents overconfidence
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure 2 (b)). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure 1 (a). Indeed, $F(N,\hat{p}(y|x))$ is smaller for easier examples (Figure 13, right), effectively regularizing their contribution. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section C.1.
@@
In practice, the introduction of $F$ can lead to some samples in a batch contributing minimally to the current gradient step. To maintain a stable effective batch size, we introduce a threshold $\epsilon$, and if for a given example, $(x,y)$, $F(N,\hat{p}(y|x))<\epsilon$, we replace the example with a new one.
Figure 2: (a) DCO improves on CE loss for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on MATH to produce direct answers: one with CE loss and others using $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for various $N^{\prime}$ (color-coded). Each curve shows pass@N coverage for a single fine-tuned model. Note that no $N^{\prime}$ is optimal for all $N$. The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. (See Figure 6, which highlights $N^{\prime}=256$, and  Figures 8 and 10 for Llama-3-70B and AIME24 results.) (b) DCO limits overconfidence by attenuating gradients for examples on which the model is highly confident. We plot the confidence regularization factor $F$ in Equation 4. For CE loss ($N=1$), $F=1$ regardless of confidence $\hat{p}(y|x)$. For $N>1$, $F$ decreases with $\hat{p}(y|x)$ and drops to zero around $1/N$—once confidence on an example reaches $O(1/N)$, $F$ vanishes, preventing it from further increasing. (c) Inverse confidence $N_{\text{eff}}$ at training time controls search-tree exploration shapes at test-time. We provide schematic proof search trees for models trained with DCOstep for small (left) and large (right) $N_{\mathrm{eff}}$. Circles are proof states; edges are tactics; solid edges are explored, while dashed edges are not. Small $N_{\mathrm{eff}}$ yields a narrow search tree that may miss correct proofs (red, dashed), while larger $N_{\mathrm{eff}}$ expands the tree to include the correct proof. However, if the tree becomes too wide ($N_{\mathrm{eff}}^{k}\gg N$ for a $k$-step proof), sampling the correct path becomes unlikely and pass@N coverage decreases as $\mathcal{C}^{N}\approx N/N_{\mathrm{eff}}^{k}$. Thus $N_{\mathrm{eff}}$, chosen at training time, is a powerful knob to control the tradeoff between exploitation and exploration at test time.
@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
We next test whether DCO can rescue test-time scaling by preventing model overconfidence. We first perform experiments on the MATH dataset in this section, and then in the next section we perform experiments on the LeanDojo automated theorem proving benchmark [42]. We fine-tune Llama-3-8B base model for 4 epochs on the MATH training set using DCO for $N^{\prime}=256$ and confirm that the model is much less confident in its greedy completion than when trained with CE loss after multiple epochs (compare Figure 1 (a) and (c)). Moreover, training with DCO at larger $N$ yields lower model greedy confidences at the end of training (Figure 1 (d)), consistent with Lemma 4.2.
We next assess pass@N test performance as a function of $N$ for models trained by minimizing $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for different values of $N^{\prime}$ (Figure 2 (a)). For any given $N$, there is an optimal $N^{\prime}$ for the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ that maximizes pass@N test coverage, yielding a Pareto optimal performance frontier (black curve) that is achieved when $N^{\prime}$ is close to $N$. In particular the model trained with CE loss (equivalent to pass@1 maximization at training) performs poorly relative to the Pareto frontier at large N (red curve below black at large $N$). Conversely, models trained with DCO at large $N^{\prime}$ perform poorly relative to the Pareto frontier for small $N$ (green curves below black at small $N$).
Together these results indicate that the alignment of the training loss $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ with the test time strategy pass@N, with $N^{\prime}$ close to $N$, is crucial for obtaining Pareto optimal performance. Moreover, these results once again confirm the tradeoff between exploration and exploitation, with good performance using pass@N test-time strategies requiring high (low) exploration with low (high) confidence at large (small) $N$.
@@ 5.3 Improved theorem proving via ensembled tree search through a modified step-wise DCO
To further test our method in realistic settings with a verifier, we conduct experiments in theorem proving using an interactive proof assistant on the LeanDojo benchmark [42] extracted from the math library of LEAN4 [43]. In this task, at each step $i$ of the proof, the model is prompted with the current proof state $x[i]$, and it outputs a proof tactic $y[i]$ with a trainable model probability $\hat{p}(y[i]|x[i])$. The proof assistant takes the sampled tactic $y[i]$, verifies whether it is a valid tactic, and if so, returns the next proof state $x[i+1]$ after the tactic $y[i]$ is applied. At test time this interactive process between the model and the assistant continues until either: (1) it terminates in an invalid tactic; (2) hits a maximal allowed search depth set by computational constraints; or (3) terminates in a successful proof of the initial proof goal as verified by the proof assistant.
@@
In our experiments, we use LEAN4 [44] as the formal proof language. We start from Qwen2.5-Math-1.5B [26], and fine-tune it on the LeanDojo benchmark [42]. In contrast to solving math problems above where the model first autoregressively generates an entire answer $y$ according to $\hat{p}(y|x)$ and then the entire $y$ is verified for correctness, in theorem proving we must obtain correct tactics $y[i]$, as verified by the proof assistant, at every proof step $i$. A straightforward application of DCO fine-tuning in this setting then involves replacing the model confidence $\hat{p}(y|x)$ in  Equation 3 on a single answer $y$, with the model confidence on an entire successful, complete $k$-step proof $\hat{p}(y[0],...,y[k-1]|x[0])=\prod_{i=0}^{k-1}\hat{p}(y[i]|x[i])$. Because of the chain rule, the DCO gradient on a single proof has the same form as Equation 4 but with $F(N,\hat{p}(y|x))$ replaced with $F(N,\prod_{i=0}^{k-1}\hat{p}\left(y[i]|x[i]\right))$. We naively applied this DCO fine-tuning method with $N^{\prime}=4$k and evaluated the resulting model, as well as a base model trained with CE loss, on MiniF2F, using pass@4k. CE loss achieves a proof success rate of 37.4%, while fine-tuning with DCO achieves a 38.7% success rate. Thus, a naive application of DCO to theorem proving by matching the parameter $N^{\prime}$ in DCO to the parameter $N$ in pass@N achieves only a modest improvement.
@@
In theorem proving, since every single intermediate proof step tactic $y[i]$ must be valid, it is natural to consider a step-wise generalization of DCO. For example, at each step $i$, the model chooses a tactic $y[i]$ with confidence $\hat{p}(y[i]|x[i])$. Our proposed step-wise DCO optimizes this single-step confidence according to Equation 4 as before, except now the step-wise confidence regularizer is given by $F(N_{\text{eff}},\hat{p}(y[i]|x[i]))$. Here one can think of $N_{\text{eff}}$ as a step-wise DCO hyperparameter that controls the exploration width at every proof step. In essence, the confidence regularizer prevents the confidence of any chosen tactic from becoming much higher than $1/N_{\text{eff}}$. Since the sum of the confidences over all tactics at each step must be $1$, this means that at test time, model exploration at each step corresponds to searching on a search tree in which each proof state $x[i]$ allows the exploration of approximately only $N_{\text{eff}}$ tactics. Thus using $N_{\text{eff}}$ in step-wise DCO during fine-tuning selects the approximate branching factor $N_{\text{eff}}$ of the proof search tree at test time (see  Figure 2 (c)).
Table 2: Proof success rates on Mathlib and MiniF2F using pass@4k. DCOstep outperforms CE loss ($N_{\mathrm{eff}}=1$) on both datasets for well-chosen $N_{\mathrm{eff}}$. Ensembling models trained with different $N_{\mathrm{eff}}$ leverages complementary strengths, with substantial gains over CE at matched test-time compute.
$N_{\mathrm{eff}}$ 1 (CE loss) 4 8 16 32 Ensemble of $N_{\mathrm{eff}}=1$
all $N_{\mathrm{eff}}$’s (5x test compute)
Mathlib 55.6% 56.4% 56.1% 56.5% 55.8% 62.2% 57.0%
MiniF2F 37.4% 39.0% 39.5% 37.0% 37.4% 43.6% 39.5%
@@
In particular, a valid (partial) proof of length $k$ will have probability of order $N_{\text{eff}}^{-k}$. Thus small $N_{\text{eff}}$ limits the exploration at test time to a tree with small branching factor, but allows longer proofs with larger numbers of steps $k$ to have higher sampling probability. This limited width search strategy should work well at test time using pass@N as long as two conditions hold: (1) a successful proof of length $k$ is present in the search tree of branching factor $N_{\text{eff}}$; and (2) the $N$ in pass@N at test time is large enough that this proof of probability $O(N_{\text{eff}}^{-k})$ is selected with high probability in $N$ passes (which starts to occur as soon as $N\gtrsim N_{\text{eff}}^{k}$). Conversely, larger $N_{\text{eff}}$ allows for more exploration at test time using a wider search tree of larger branching factor, but it makes finding longer proofs with large $k$ harder, since the approximate success condition for pass@N of $N>N_{\text{eff}}^{k}$ is harder to satisfy at fixed $N$ and large $N_{\text{eff}}$ and $k$. However, this wide exploration strategy can work well for short proofs with small $k$ if a successful short proof does not lie in the limited width search tree obtained at small $N_{\text{eff}}$, but does lie in a wider search tree at larger $N_{\text{eff}}$ (Figure 2, (c) right).
@@
The mean and median lengths of the proofs in the training set are 4.2 and 2, respectively. Thus we explore fine-tuning with the step-wise DCO algorithm, DCOstep, with $N_{\text{eff}}$ ranging from $1$ (CE loss) to $N_{\text{eff}}=32$. We evaluate pass@4k performance on the Mathlib and MiniF2F test sets and find that larger $N_{\mathrm{eff}}$ improves accuracy over the CE baseline (Table 2). Interestingly, the optimal $N_{\text{eff}}$ increases with more passes at test time (Table 6), similar to the Pareto frontier in Figure 2 (a). Since different choices of $N_{\text{eff}}$ correspond to different search strategies at test time—larger $N_{\text{eff}}$ favors short proofs in wide trees, smaller $N_{\text{eff}}$ favors long proofs in narrow trees— we hypothesized that ensembling these methods could significantly boost performance. This was indeed the case (ensemble entry in Table 2). The ensemble strategy samples $5\times 4$k times, so a proper baseline is CE loss with the same test-time compute of pass@N with $N=20$k. The ensemble strategy outperforms this stringent baseline with significant excesses of $5.2\%$ on Mathlib and $4.1\%$ on MiniF2F.
@@
These results indicate that varying $N_{\text{eff}}$ in DCOstep at training time allows a diversity of tree search strategies trading depth and breadth that can be exploited by scaling test-time compute via pass@N.
@@
5.4 Approximate DCO also improves performance with chain-of-thought reasoning
Figure 3: (a, b) Overconfidence persists in CoT fine-tuning with CE loss. We fine-tune a Llama-3-8B base model on CoT traces in the MATH dataset and plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. (a) For CE loss, the model becomes more confident in its most likely answers as training progresses—the confidence distribution shifts to the right. This effect is milder than in the direct answer setting (Figure 1). (b) DCOa limits this shift of the confidence distribution over training. (c, d) Overconfidence is also present in GRPO fine-tuning and can result in a model that is overconfident and wrong. (c) The model trained with GRPO assigns progressively larger confidences to $\hat{p}(y^{\mathrm{mode}}|x)$ over the course of training. (d) At the end of training, only a small portion of the model’s highly confident completions are correct, which will harm the model’s pass@N performance when scaling up $N$.